# SO-Bench: A Structural Output Evaluation of Multimodal LLMs

### v1

@@ 1 Introduction
While structured output from text-only inputs (yang2025structeval; tang2023struc; gu2024structext; geng2025jsonschemabench) has received growing attention, the multimodal domain, especially the visual structured output capability of MLLMs, remains largely underexplored. Given visual inputs (e.g., natural images, screens, charts) and customized schemas (often specified by downstream applications), visual structured output involves extracting information from images, and aligning them with hierarchical schemas and user instructions. For instance, adding the product labels shown on a UI screen to the shopping cart, or saving the information from the menu to the notes apps, illustrated in Figure 1. Prior visual-language works such as Pix2Struct (lee2023pix2struct), Image2Struct (roberts2024image2struct), and IR3D-Bench (liu2025ir3d) have made progress toward structured visual understanding, but they typically emphasize text captioning or semantic parsing with limited domains rather than rigorous customized schema-level adherence with diverse real-world scenarios. Most relevant to visual structured output is key information extraction (KIE) in document parsing and OCR understanding (yang2024cc; ouyang2025omnidocbench), where a MLLM is tasked to extract information for pre-defined keywords. However, KIE focuses on semantic parsing with simple input fields, ignoring the diversity of real-world schemas with nested, complex structure. As a result, there is no study that systematically quantifies how well MLLMs can produce structured, schema-conformant outputs grounded in visual evidence.
@@
Based on SO-Bench, we conduct comprehensive benchmarking experiments across a wide range of MLLMs of different scales, including open-sourced and proprietary models. Our results reveal that small models (e.g., with 3B and 7B scales) exhibit substantial gaps in structured generation and schema compliance compared to their proprietary counterparts. Though the strongest frontier models (e.g., GPT-5 and Gemini 2.5-Pro) show good performance in schema following with over $95\%$ accuracy, there is a large room of improvements for predicting fully correct structured outputs, with only up to $17.1\%$ accuracy. Ablation studies further illustrate that a model’s visual structured output ability highly correlates with its competence in visual instruction following, agentic tool invocation and parameter filling, and general visual knowledge.
To study how training impacts models’ visual structured output capabilities, we further construct a large-scale training set using the same data labeling pipeline, and train MLLMs with supervised finetuning (SFT) and Reinforcement Learning with Verfiable Rewards (RLVR). Experimental results show that both SFT and RLVR enhances schema validation and field matching accuracy up to $20\%$ and $13\%$, respectively, underscoring the importance of targeted supervision for improving structured reasoning.
@@ 2 Related Works
Structured Output Benchmarks. Several benchmarks have been recently proposed to quantify a LLM’s capability to generate structured output with text-only inputs. For example, StructEval yang2025structeval benchmarks text-to-structure generation across formats like JSON, YAML, and HTML, emphasizing format fidelity. JSONSchemaBench geng2025jsonschemabench extends this by testing constrained decoding on thousands of real-world JSON schemas, measuring coverage and quality across diverse constrained decoding frameworks. StructBench tang2023struc auto-generates structure-rich text tasks such as table-to-JSON or LaTeX conversion. In contrast to text-only applications, visual structured output focuses on extracting image information following the pre-defined schemas. Pix2Struct lee2023pix2struct pretrains an image-to-text model that parses screenshots to HTML, improving layout understanding. Image2Struct roberts2024image2struct targets structured extraction from rendered data (webpages, formulas, music scores), yet only focuses on digital images with limited domains. IR3D-Bench liu2025ir3d tests structured scene reconstruction from pure synthetic 3D renderings. Overall, existing studies either omit image inputs or cover narrow visual domains, leaving the broader capability of MLLMs in generating diverse, schema-grounded outputs unexplored.
@@
SO-Bench also relates to agentic tool-use evaluations that quantify structured reasoning and function-call accuracy. The Berkeley Function Calling Leaderboard (BFCL) (patilberkeley) and Tau-Bench (yao2024tau) assess agents’ ability to produce executable structured calls and manage multi-step tool interactions, while ToolVQA yin2025toolvqa extends this concept to visual question answering with external tools. Unlike these benchmarks, which focus on textual tool APIs or with limited number of tools (e.g., $14$ tools in GTA (wang2024gta)), SO-Bench evaluates multimodal structured generation grounded in visual inputs with large-scale diverse JSON schemas, bridging perception and structured reasoning.
@@
Visual structured output connects to document understanding and OCR-rich tasks, as it requires a model to correctly parse textual information from images. Benchmarks such as OCRBenchV2 (fu2024ocrbench), CC-OCR (yang2024cc), and OmniDocBench (ouyang2025omnidocbench) focus on extracting text and layout from scanned or synthetic documents, often formulating key information extraction (KIE) tasks with simple predefined templates in one-layer dictionary. In contrast, SO-Bench focuses on distinct, realistic JSON schemas with nested structures, aligned with application-driven requirements.
@@ 3.1 Overview
We formulate the visual structured output problem as follows: given an input image $I$, a pre-defined JSON schema $S$, and a user instruction $X$, a model is expected to generate structured output $Y$ that both (1) conforms syntactically to $S$ and (2) semantically reflects the information extracted from $I$ and $X$. For a multimodal LLM, the structured output is generated auto-regressively with the predicted probability $p(Y|I,X,S)$. A JSON schema is a nested dictionary specifying keys, data types, and object hierarchy. A user instruction provides guidance on the information to be extracted from the input images, which can be precise and descriptive, or concise and ambiguous, for example, “Help me save this poster".
The diversity of our benchmark originates from images (domain coverage and visual representations) and JSON schemas (e.g., schema complexity, field type variety), introduced in Section 3.2. The main challenge for building such a benchmark is how to associate an image with the representative JSON schema, and generate its structured output response accurately and efficiently. We develop a multi-stage autolabeling pipeline with human in-the-loop to tackle this challenge, detailed in Section 3.3.
@@ 3.2 Data Collection
To ensure broad visual coverage, SO-Bench aggregates image data from ten recent public datasets spanning four domains: (1) UI (mobile, web, and desktop screens), (2) natural images, (3) documents, and (4) charts. These domains jointly cover a wide range of real-world structured extraction scenarios, such as menu parsing, receipt understanding, poster collection, UI layout analysis, document parsing, and chart interpretation. We prioritize English-language data and selectively down-sample test sets to maximize visual and semantic diversity. Figure 3(a) summarizes the data composition and domain distribution. Regarding JSON schema collection, we curated a large and diverse schema repository covering both real-world and synthetic cases. The repository includes a subset of the real-world JSON schemas from JsonSchemaBench (geng2025jsonschemabench), which spans domains such as function signatures, service APIs, and system configurations, we keep 66K unique schemas related to visual applications. We additionally generate $500$ synthetic JSON schemas by prompting a proprietary model for specific use cases (e.g., reading receipt, parsing nutrition labels). This collection serves as a JSON schema databank from which image–schema pairs are constructed.
@@
To enable efficient schema–image retrieval and matching (detailed in Section 3.3), we embed both images and JSON schemas using an off-the-shelf CLIP model (radford2021learning). For an Image $I$, we first use a proprietary model (e.g., GPT-5 or Gemini-2.5-Pro) to generate its dense caption $T$, and then employ the CLIP’s image and text encoders to extract visual and textual embeddings, denoted by $E_{I}$ and $E_{T}$, respectively. As for a JSON schema $S$, we prompt with another proprietary model to summarize its fields, purpose, and example use cases with a canonical name (e.g., ReceiptExtractor, MenuReader). These textual descriptions are embedded with the CLIP text encoder, denoted by $E_{S}$.
@@ 3.3.3 Progressive Response Generation and Refinement
For each triplet of image, schema, and user intent, we generate the structured output using our response generation pipeline. When available, auxiliary signals such as OCR text, ground truth values, layout metadata, or HTML structures (for UI images) are provided as guidance to improve fidelity. The generated JSON outputs are validated against schema constraints to ensure syntactic correctness and logical completeness. Besides, we introduce a hybrid automated and manual review loop. A critic–refiner workflow, powered by LLM-based validators, examines each structured output for schema validity and semantic consistency and summarizes suggestions of improvements. Invalid or sub-optimal outputs are regenerated up to three times considering the feedbacks from the critic model. During this process, human experts conduct manual inspections and provide feedbacks for the next round output refinement.
@@ 3.5 Evaluation Metrics
We used abstract syntax tree (AST) evaluation method following BFCL (patilberkeley). In particular, given the structured output from the model and ground truth in dict format, we iteratively compare all keys in the dict and their corresponding values, and recursively compare all the nested dictionaries. For all primitive types (int, float, str, list), we used exact match by default. Since not all information in the ground truth is directly relevant to the user intent, and some visual information can be hard to extract exactly (e.g., blurry, truncated images), we further define two other matching strategies for primitive types: 1) we allow fuzzy match if the ground truth information does not include explicit texts shown in the image (e.g., exact data value in a line plot). Here, we adopt normalized edit distance to compare strings and relative error to compare int and float. 2) We also label ignore for certain fields, if the key is not a required field in the schema definition and it is not directly related to the user query (e.g., the style of an UI screenshot when asking about content). Here, we simply assign a matching score of $1$, if the fuzzy match score is above a pre-defined threshold and skip comparisons for ignored fields. To enable such fine-grained evaluation, we also add an evaluation label generation step in our pipeline, where we provide the image, schema, user intent and ground-truth response (after final refinement), and prompt a proprietary MLLM to generate a matching type for each primitive field (e.g., one of {exact, fuzzy, ignore}).
We report the following three metrics in our experiments. (1) Schema Validation Accuracy: the percentage of outputs that are valid w.r.t the given schema definition. (2) Field Matching Accuracy: Let $\mathcal{F}(D)$ denote all fields in nested dictionary $D$ (including both intermediate nested structures and leaf fields). For ground truth $G^{(k)}$ and output $O^{(k)}$ across $N$ examples:
@@
where $\text{Match}(f,f^{\prime})$ returns true for exact/fuzzy matches, with nested structures matching when all their subfields match. (3) Full Structure Matching Accuracy:
@@ 4.1 Main Results
Table 1 compares the structured output performance across several state-of-the-art open-source and large proprietary models. We report three metrics, namely, schema validation accuracy, field matching accuracy (exact and fuzzy), and full structure matching accuracy (exact and fuzzy), as introduced in Section 3.5. Among all models, Gemini-2.5-Pro achieves top results across all metrics, with a schema validation accuracy reaching nearly $98\%$, demonstrating strong schema adherence capabilities. GPT-5 also shows a clear improvement over its predecessor GPT-4o with over $15\%$ higher schema validation accuracy. However, all models exhibit relatively lower performance in field matching. For field match (fuzzy), only Gemini-2.5-Pro achieves accuracy higher than $70\%$, and no model achieves $18\%$ on Full Match (fuzzy), indicating substantial room for improvement in producing fully compliant structured outputs. Upon further inspection of errors, we also found that the large gap between field match and full match accuracies are often due to the model predicting semantically correct answers for certain fields, which is beyond what fuzzy match can capture. We leave semantic matching methods out from our metrics for simplicity and reproducibility, and we think developing more flexible matching functions could be an interesting future work direction.
@@ 4.2.3 Inference with Structured Output API
Schema Val. Field Match Schema Val. Field Match
@@
Table 3: Results from structured output APIs and instruction following prompts.
@@ 4.3 Training Experiments
To study how training impacts the performance of models on our SO-Bench, we further constructed a large-scale training set using the same pipeline as described in Section 3.3, except for the human verification part. In particular, we source images from the training set of HierText (long2022towards), AriaUI (yang2025aria) and COYO (kakaobrain2022coyo-700m), and paired them with either our JSON schema repository or syntehtic JSON schemas . We then proceed with persona-based user intent generation and response generation followed by final model-based refinement. In total, we collected 114K examples for training. We then proceed with SFT and RL experiments on this data to understand its impact. More details on training experiments are in Appendix B.
@@ B.1 Training Dataset
This section provides additional details on the training dataset used in our study (Section 4.3). For supervised fine-tuning, we construct a training corpus sourced from the training sets of HierText (long2022towards), AriaUI (yang2025aria), and COYO (kakaobrain2022coyo-700m). We balance both in-domain vs. out-of-domain data and real vs. synthetic JSON schemas. HierText serves as the in-domain dataset because its test split is used to build part of the SO-Bench evaluation set. COYO and AriaUI, in contrast, provide out-of-domain training data. For the full HierText training set (10K images), we generate one version with synthetic schemas and another with real JSON schemas. For AriaUI, we downsample to 1010K images, drawn from desktop, mobile, and web UI screens. Similar to HierText split, we generate one example with synthetic schema and another with real JSON schema for every image. Since COYO is extremely large (over 747M image–text pairs), we first categorize and then randomly downsample it to 9595K images, consisting of: 2020K infographics, 1010K handwritten images, 4040K scene-text images, 55K maps, and 2020K tables. Among these, $80\%$ of the images are paired with real JSON schemas, while $20\%$ use synthetic schemas. Finally, we filter out noisy samples with low label quality, ending up with 114114K training samples.

### v2

@@ 1 Introduction
While structured output from text-only inputs (yang2025structeval; tang2023struc; gu2024structext; geng2025jsonschemabench) has received growing attention, the multimodal domain, especially the visual structured output capability of MLLMs, remains largely underexplored. Given visual inputs (e.g., natural images, screens, charts) and customized schemas (often specified by downstream applications), visual structured output involves extracting information from images, and aligning them with hierarchical schemas and user instructions. For instance, adding the product labels shown on a UI screen to the shopping cart, or saving the information from the menu to the notes apps, illustrated in Figure 1. Prior visual-language works such as Pix2Struct (lee2023pix2struct), Image2Struct (roberts2024image2struct), and IR3D-Bench (liu2025ir3d) have made progress toward structured visual understanding, but they typically emphasize text captioning or semantic parsing with limited domains rather than rigorous customized schema-level adherence with diverse real-world scenarios. Most relevant to visual structured output is key information extraction (KIE) in document parsing and OCR understanding (yang2024cc; ouyang2025omnidocbench), where a MLLM is tasked to extract information for pre-defined keywords. However, KIE focuses on semantic parsing with simple input fields, ignoring the diversity of real-world schemas with nested, complex structure. As a result, there is no study that systematically quantifies how well MLLMs can produce structured, schema-compliant outputs grounded in visual evidence.
@@
Based on SO-Bench, we conduct comprehensive benchmarking experiments across a wide range of MLLMs of different scales, including open-sourced and proprietary models. Our results reveal that small models (e.g., with 3B and 7B scales) exhibit substantial gaps in structured generation and schema compliance compared to their proprietary counterparts. Though the strongest frontier models (e.g., GPT-5 and Gemini 2.5-Pro) show good performance in schema following with over $95\%$ accuracy, there is a large room of improvements for predicting fully correct structured outputs, with only up to nearly $19\%$ accuracy (fuzzy match). Ablation studies further illustrate that a model’s visual structured output ability highly correlates with its competence in visual instruction following, agentic tool invocation and parameter filling, and general visual knowledge.
To study how training impacts models’ visual structured output capabilities, we further construct a large-scale training set at the post-training stage using the same data labeling pipeline, and train MLLMs with supervised finetuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). Experimental results show that both SFT and RLVR enhances schema validation and field matching accuracy up to $20\%$ and $13\%$, respectively, underscoring the importance of targeted supervision for improving structured reasoning.
@@ 2 Related Works
Structured Output Benchmarks. Several benchmarks have been recently proposed to quantify a LLM’s capability to generate structured output with text-only inputs. For example, StructEval yang2025structeval benchmarks text-to-structure generation across formats like JSON, YAML, and HTML, emphasizing format fidelity. JSONSchemaBench geng2025jsonschemabench extends this by testing constrained decoding on thousands of real-world JSON schemas, measuring coverage and quality across diverse constrained decoding frameworks. StructBench tang2023struc auto-generates structure-rich text tasks such as table-to-JSON or LaTeX conversion. In contrast to text-only applications, visual structured output focuses on extracting image information following the pre-defined schemas. Pix2Struct lee2023pix2struct pretrains an image-to-text model that parses screenshots to HTML, improving layout understanding. Image2Struct roberts2024image2struct targets structured extraction from rendered data (webpages, formulas, music scores), yet only focuses on digital images with limited domains. IR3D-Bench liu2025ir3d tests structured scene reconstruction from pure synthetic 3D renderings. Overall, existing studies either omit image inputs or cover narrow visual domains, leaving the broader capability of MLLMs in generating diverse, schema-grounded outputs under-explored.
@@
SO-Bench also relates to agentic tool-use evaluations that quantify structured reasoning and function-call accuracy. The Berkeley Function Calling Leaderboard (BFCL) patilberkeley and Tau-Bench yao2024tau assess agents’ ability to produce executable structured calls and manage multi-step tool interactions, while ToolVQA yin2025toolvqa extends this concept to visual question answering with external tools. Unlike these benchmarks, which focus on textual tool APIs or with limited number of tools (e.g., $14$ tools in GTA (wang2024gta)), SO-Bench evaluates multimodal structured generation grounded in visual inputs with large-scale diverse JSON schemas, bridging perception and structured reasoning.
@@
Visual structured output connects to document understanding and OCR-rich tasks, as it requires a model to correctly parse textual information from images. Benchmarks such as OCRBenchV2 (fu2024ocrbench), CC-OCR (yang2024cc), and OmniDocBench (ouyang2025omnidocbench) focus on extracting text and layout from scanned or synthetic documents, often formulating key information extraction (KIE) tasks with simple predefined templates in one-layer dictionary. In contrast, SO-Bench focuses on distinct, realistic JSON schemas with nested structures, aligned with application-driven agentic tool-use requirements.
@@ 3.1 Overview
We formulate the visual structured output problem as follows: given an input image $I$, a pre-defined JSON schema $S$, and a user instruction $X$, a model is expected to generate structured output $Y$ that both (1) conforms syntactically to $S$ and (2) semantically reflects the information extracted from $I$ and $X$. For a multimodal LLM, the structured output is generated auto-regressively with the predicted probability $p(Y|I,X,S)$. A JSON schema is a nested dictionary specifying keys, data types, and object hierarchy following the standard (jsonschema). A user instruction provides guidance on the information to be extracted from the input images, which can be precise and descriptive, or concise and ambiguous, for example, “Help me save this poster”.
The benchmark should be diverse in order to test the general visual structured output capabilities of MLLMs. The diversity originates from images (domain coverage and visual representations) and JSON schemas (e.g., schema complexity, field type variety), which will be introduced in Section 3.2. The main challenge for building such a benchmark is how to associate an image with the representative JSON schema, and generate its structured output response accurately and efficiently. To tackle this challenge, we develop a multi-stage auto-labeling pipeline with human in-the-loop detailed in Section 3.3.
@@ 3.2 Data Collection
To ensure broad visual coverage, SO-Bench aggregates image data from ten recent public datasets spanning four domains with different image resolutions and aspect ratios: (1) UI (mobile, web, and desktop screens), (2) natural images, (3) documents, and (4) charts. These domains jointly cover a wide range of real-world structured extraction scenarios, such as menu parsing, receipt understanding, poster collection, UI layout analysis, document parsing, and chart interpretation. We prioritize English-language data and selectively down-sample test sets to maximize visual and semantic diversity. Figure 3(a) summarizes the data composition and domain distribution. Regarding JSON schema collection, we curated a large and diverse schema repository covering both real-world and synthetic cases. The repository includes a subset of the real-world JSON schemas from JsonSchemaBench (geng2025jsonschemabench), which spans domains such as function signatures, service APIs, and system configurations. Out of 1111K full schema repository, we keep 66K unique schemas related to visual applications. We additionally generate $500$ synthetic JSON schemas by prompting a proprietary model for specific use cases (e.g., reading receipt, parsing nutrition labels). This collection serves as a JSON schema databank from which image–schema pairs are constructed.
@@
To enable efficient schema–image retrieval and matching (detailed in Section 3.3), we embed both images and JSON schemas using an off-the-shelf CLIP model (radford2021learning). For an Image $I$, we first use a proprietary model (e.g., GPT-5 or Gemini-2.5-Pro) to generate its caption $T$, and then employ the CLIP’s image and text encoders to extract visual and textual embeddings, denoted by $E_{I}$ and $E_{T}$, respectively. As for a JSON schema $S$, we prompt with another proprietary model to summarize its fields, purpose, and example use cases with a canonical name (e.g., ReceiptExtractor, MenuReader). These textual descriptions are embedded with the CLIP text encoder, denoted by $E_{S}$.
@@ 3.3.1 Image-Schema Association
Note that the semantic search approach described above could still end up with less-relevant image-schema associations. We performed human quality check and filter out those samples before proceeding to the next data generation stage.
@@ 3.3.3 Progressive Response Generation and Refinement
For each triplet of image, schema, and user intent, we generate the structured output using our response generation pipeline. When available, auxiliary signals such as OCR texts, layout metadata, or HTML structures (for UI images) provided by the original benchmark datasets are used as guidance to improve fidelity. The generated JSON outputs are validated against schema constraints to ensure syntactic correctness and logical completeness. Besides, we introduce a hybrid automated and manual review loop. A critic–refiner workflow, powered by LLM-based validators, examines each structured output for schema validity and semantic consistency and summarizes suggestions of improvements. Invalid or sub-optimal outputs are regenerated up to three times considering the feedback from the critic model. During this process, human experts conduct manual inspections and provide feedback for the output refinement.
@@ 3.5 Evaluation Metrics
We used abstract syntax tree (AST) evaluation method following BFCL (patilberkeley). In particular, given the structured output from the model and ground truth in dict format, we iteratively compare all keys in the dict and their corresponding values, and recursively compare all the nested dictionaries. For all primitive types (int, float, str, list), we used exact match by default. Since not all information in the ground truth is directly relevant to the user intent, and some visual information can be hard to extract exactly (e.g., blurry, truncated images), we further define two other matching strategies for primitive types: 1) we allow fuzzy match if the ground truth information does not include explicit texts shown in the image (e.g., exact data value in a line plot). Here, we adopt normalized edit distance to compare strings and relative error to compare int and float. 2) We also label ignore for certain fields, if the key is not a required field in the schema definition and it is not directly related to the user query (e.g., the style of an UI screenshot when asking about content). Here, we simply assign a matching score of $1$, if the fuzzy match score is above a pre-defined threshold and skip comparisons for ignored fields. To enable such fine-grained evaluation, we also add an evaluation label generation step in our pipeline, where we provide the image, schema, user intent and ground-truth response (after final refinement), and prompt a proprietary MLLM to generate a matching type for each primitive field (e.g., one of {exact, fuzzy, ignore}). The pseudo code of our evaluation function is presented in Algorithm 1.
We report the following three metrics in our experiments. (1) Schema Validation Accuracy: the percentage of outputs that are valid w.r.t the given schema definition. (2) Field Matching Accuracy (FMA): Let $\mathcal{F}(D)$ denote all fields in nested dictionary $D$ (including both intermediate nested structures and leaf fields). For ground truth $G^{(k)}$ and output $O^{(k)}$ across $N$ examples:
@@
where $\text{Match}(f,f^{\prime})$ returns true for exact/fuzzy matches, with nested structures matching when all their subfields match. (3) Full Structure Matching Accuracy (FSMA):
@@ 4.1 Main Results
Table 1 compares the structured output performance across several state-of-the-art open-source and large proprietary models. We report three metrics, namely, schema validation accuracy, field matching accuracy (exact and fuzzy), and full structure matching accuracy (exact and fuzzy), as introduced in Section 3.5. Among all models, Gemini-2.5-Pro achieves top results across all metrics, with a schema validation accuracy reaching nearly $98\%$, demonstrating strong schema adherence capabilities. GPT-5 also shows a clear improvement over its predecessor GPT-4o with over $15\%$ higher schema validation accuracy. However, all models exhibit relatively lower performance in field matching. For field match (fuzzy), only Gemini-2.5-Pro achieves accuracy higher than $70\%$, and no model achieves above $20\%$ on Full Match (fuzzy), indicating substantial room for improvement in producing fully compliant structured outputs. Upon further inspection of errors, we also found that the large gap between field match and full match accuracies are often due to the model predicting semantically correct answers for certain fields, which is beyond what fuzzy match can capture. We leave semantic matching methods out from our metrics for simplicity and reproducibility, and we think developing more flexible matching functions could be an interesting future work direction.
@@ 4.2.3 Inference with Structured Output API
Schema Val. Field Match (Fuzzy) Schema Val. Field Match (Fuzzy)
@@
Table 3: Results from structured output APIs and instruction following prompts on a subset of SO-Bench.
@@ 4.3 Training Experiments
To study how training impacts the performance of models on our SO-Bench, we further constructed a large-scale training set using the same pipeline as described in Section 3.3, except for the human verification part. In particular, we source images from the training set of HierText (long2022towards), AriaUI (yang2025aria) and COYO (kakaobrain2022coyo-700m), and paired them with either our JSON schema repository or synthetic JSON schemas . We then proceed with persona-based user intent generation and response generation followed by final model-based refinement. In total, we collected 114K examples for training. We then proceed with SFT and RL experiments on this data to understand its impact. More details on training experiments are in Appendix B.
@@ 5 Conclusion
Discussions and Limitations
We identify several limitations of our work here. First, due to the high complexity of generating structured output, we rely on frontier models to automatically generate the initial benchmark labels. We applied several strategies to improve the label quality, including 1). multi-stage auto-labeling pipeline design, 2). per-stage correction from human annotators and critic models, 3). shuffling among proprietary models (e.g. GPT-5 or Gemini-2.5-Pro) at each generation stage to reduce model biases. However, there could still be some label errors which cap a model’s performance. Secondly, as we discussed in the section 4.1, there could exist multiple correct ground truths for some examples and our current annotation only captures one of them. For example, semantically similar sentences for string fields, or certain flexible schema that allow different realization. Thus the models’ performance might be underestimated for those cases. Finally, in our evaluation metrics design, FMA and FSMA (Equation 3.3 and Equation 3.4) mainly capture the recall rates of a prediction. A model which predicts many additional fields not existed in ground truth labels could still achieve high FMA or FSMA scores. We leave better metrics design as future work.
@@ B.1 Training Dataset
This section provides additional details on the training dataset used in our study (Section 4.3). For supervised fine-tuning, we construct a training corpus sourced from the training sets of HierText (long2022towards), AriaUI (yang2025aria), and COYO (kakaobrain2022coyo-700m). We balance both in-domain vs. out-of-domain data and real vs. synthetic JSON schemas. HierText serves as the in-domain dataset because its test split is used to build part of the SO-Bench evaluation set. COYO and AriaUI, in contrast, provide out-of-domain training data. For the full HierText training set (10K images), we generate one version with synthetic schemas and another with real JSON schemas. For AriaUI, we downsample to 1010K images, drawn from desktop, mobile, and web UI screens. Similar to HierText split, we generate one example with synthetic schema and another with real JSON schema for every image. Since COYO is extremely large (over 747M image–text pairs), we first categorize and then randomly downsample it to 9595K images, consisting of: 2020K infographics, 1010K handwritten images, 4040K scene-text images, 55K maps, and 2020K tables. Among these, $80\%$ of the images are paired with real JSON schemas, while $20\%$ use synthetic schemas. Finally, we filter out noisy samples with low label quality using the critic model, ending up with 114114K training samples.